# Cosmological effect of coherent oscillation of ultralight scalar fields in a multicomponent universe

### v1

@@ Abstract
The idea that coherent oscillations of a scalar field, oscillating over a time period that is much shorter than the cosmological timescale, can exhibit cold dark matter (CDM) like behavior was previously established. In our work we first show that this equivalence between the oscillating scalar field model and the CDM sector is exact only in a flat Friedmann-Lemaitre-Robertson-Walker (FLRW) spacetime in the absence of cosmological constant and any other possible matter components in the universe when the mass of the scalar field is very large compared to the Hubble parameter. Then we show how to generalize the equivalence between the coherently oscillating scalar field model and the CDM sector in a spatially curved universe with multiple matter components. Using our general method, we will show how a coherently oscillating scalar field model can represent the CDM sector in the presence of non-minimal coupling of the CDM sector with radiation. Our method is powerful enough to work out the dynamics of gravitational collapse in a closed FLRW spacetime where the coherently oscillating scalar field model represents the CDM sector. We have, for the first time, presented a consistent method which specifies how a coherently oscillating scalar field model, where the scalar field is ultralight, acts like the CDM sector in a multicomponent universe.
@@ I Introduction
The standard model of cosmology attributes approximately 27 percent of the universe’s total energy density to dark matter, an as-yet unidentified component that interacts predominantly through gravity. This estimate is supported by the final results of the full mission of the Planck satellite Planck:2018vyg , which constrain cosmological parameters with high precision. These results are interpreted within the framework of $\Lambda$CDM cosmology, where $\Lambda$ stands for the cosmological constant and CDM specifies a cold dark matter constituent of the universe. Generally, the CDM sector is modeled by an ideal fluid with negligible pressure. There are various ways in which people have tried to model the CDM fluid, and one of the interesting ways, out of many, is related to a model of coherently oscillating scalar field as proposed by Turner in Ref. PhysRevD.28.1243 . Turner showed that for a quadratic potential for the scalar field, there appear to be two distinct timescales: the intrinsic oscillation period of the field related to $1/m$, where $m$ is the mass of the scalar field, and the cosmological time scale related to $1/H$, where $H$ is the Hubble parameter. When the scalar field mass satisfies $m\gg H$, the field undergoes rapid oscillations with a period $\sim m^{-1}$, significantly shorter than the cosmological time scale $H^{-1}$. As a result, the field completes many oscillations within a single Hubble time, and its time-averaged dynamics closely resemble the dynamics of CDM in an expanding Friedmann-Lemaitre-Robertson-Walker (FLRW) spacetime. The coherently oscillating scalar field exhibits zero average pressure, and the average energy density evolves as the inverse cube of the cosmic scale factor.
@@
Ratra PhysRevD.44.352 later extended the analysis of Turner and produced a consistent mathematical method to deal with coherently oscillating scalar fields in cosmology. Ratra consistently worked out the background cosmological dynamics of the universe, where only the scalar field was a matter component, and then used a synchronous gauge to calculate cosmological perturbations on the background model. The scalar field dark matter model has been extensively explored in the literature. Several subsequent studies Baer:2014eja ; Hidalgo:2017dfp ; Hotinli:2021vxg ; Garcia:2022vwm ; ChoeJo:2023ffp have explicitly adopted the approach introduced by Turner and Ratra. Hwang HWANG1997241 introduced an alternative approach by employing the uniform-curvature gauge to derive the perturbed equations, demonstrating its suitability for scalar field models. Importantly, although these studies use different gauges for spatial perturbations, the background evolution remains consistent across these approaches. The linear perturbation framework has also been applied to general power-law (scalar field) potentials, with the case $n=2$ treated as a special instance Cembranos:2015oya . This analysis builds upon previous work and demonstrates that the resulting solutions closely resemble the perturbative behavior of dust-like perfect fluids. These methods have been further developed to include second-order perturbations Noh:2013coa and extended to fully nonlinear analysis Noh:2017sdj , reinforcing the result that coherently oscillating scalar fields (or axion-like fields) exhibit cold dark matter-like behavior. Additionally, extensions of the scalar field dark matter framework have considered corrections to the standard behavior McDonald:1993ky ; Cembranos:2018ulm , including higher-order interactions or modifications to the potential. Other studies have incorporated self-interacting terms within the $\phi^{2}$ potential in the context of interacting dark sector models Aboubrahim:2024spa ; vandeBruck:2022xbk , further broadening the scope and applicability of scalar field dark matter scenarios.
@@
Investigations into the dynamical and gravitational instabilities associated with oscillating scalar fields have also played a crucial role in assessing their cosmological viability. A detailed analysis by Johnson:2008se demonstrated that scalar fields with a negative equation-of-state parameter possess a large-scale dynamical instability for the growth of perturbations. This instability does not occur when the scalar field exhibits a vanishing equation-of-state parameter, making such configurations unsuitable for explaining cosmic acceleration but remaining viable as dark-matter candidates. The growth of cosmological perturbations in scalar field dark matter models has been further examined in works like Alcubierre:2015ipa ; Marsh:2010wq , confirming their relevance for structure formation scenarios. Furthermore, extensive numerical comparisons with the standard cold dark matter (CDM) paradigm, such as those in Torres:2014bpa , have shown that scalar field and CDM perturbations evolve nearly identically in the linear regime. The primary difference lies in the presence of a characteristic cutoff in the power spectrum of the density perturbation, introduced by the scalar field dynamics. To improve the accuracy of these models, additional efforts have been made to refine the effective fluid approximation for coherently oscillating scalar field dark matter Passaglia:2022bcr , thereby enhancing their consistency with observational data.
In the context of interacting models, several studies have adopted the final results from the scalar field dark matter framework developed by Ratra and others, applying them directly to systems involving additional components or interactions HWANG20091 ; Mishra:2017ehw ; Park:2012ru ; Cardenas:2020nny ; Brax:2019fzb ; Urena-Lopez:2015gur . In these works, the oscillating scalar field is treated as dark matter without explicitly deriving the constraint on the initial condition of the oscillating ultralight scalar field. It is worth noting that the term ultralight scalar field was not originally used in the foundational works; it became common in the literature only later, and this has been further discussed in the following paragraph. The authors of Ref. Mishra:2017ehw hinted at the fine tuning of the initial value of the ultralight scalar field. The equivalence between a coherently oscillating ultralight scalar field model (oscillating in a time period much shorter than the cosmological time scale) in a spatially flat FLRW spacetime and the CDM sector holds exactly when the universe has only one matter component: the ultralight scalar field. The presence of multiple components, with minimal or non-minimal coupling between them, can significantly alter the equivalence of the two systems. The presence of other components may prevent the equivalence of the two sectors, and one may not have a proper dust-dominated phase, potentially invalidating the very basis of the equivalence. It therefore remains unclear whether the scalar field sector will continue to represent the dark matter sector in more general settings. Modifications to the existing framework are therefore necessary that consistently capture oscillating scalar field dynamics in a multicomponent universe. In our work, we develop a general method using which one can maintain the equivalence between the coherently oscillating ultralight scalar field model and the CDM sector even in a multicomponent universe. Our method works for CDM models even in the end phase of radiation domination or in the dark energy-dominated phase. The calculations presented in this paper, for the first time, make the coherently oscillating ultralight scalar field model of CDM very general and applicable to various phases of cosmological dynamics. After formulating a theoretical framework, we also validate its applicability through numerical analysis. For this analysis, we select a mass range of the ultralight scalar field that is cosmologically significant and widely recognized in the study of scalar field dark matter.
@@
The phrase mass of the scalar field frequently appears in the literature, and it is worth clarifying its intended meaning. Typically, it refers to the mass of the quantum of the scalar field—that is, the associated particle. This identification becomes natural and direct when working in natural units, where physical quantities like mass and energy are often interchangeable. In this work, however, we adopt the geometrized unit system in which $18\pi G=c=$ 1, where $G$ denotes Newton’s gravitational constant and $c$ is the speed of light. Under this convention, both the scalar field mass $m$ and the Hubble parameter $H$ are expressed in units of inverse length, specifically ${\rm cm}^{-1}$. Unlike in natural units, where such quantities would naturally carry dimensions of mass or energy—our chosen system does not assign them such interpretations directly. Furthermore, we do not quantize the scalar field in this work; our analysis is entirely classical. Nevertheless, for ease of communication and consistency with standard terminology in the field, we continue to refer to $m$ as the “mass” of the dark matter particle—while keeping in mind that, in our unit system, it has dimensions of inverse length ${\rm cm}^{-1}$.
To demonstrate the robustness of our general approach in the presence of other cosmological components, we extend our analysis by incorporating dark energy in the form of the traditional cosmological constant $\Lambda$. It is seen that our method can incorporate the cosmological constant. The standard $\Lambda$CDM paradigm faces several theoretical and observational challenges DelPopolo:2016emo ; Perivolaropoulos:2021jda . In response to these issues, dynamical dark energy models based on scalar fields have been widely proposed. Among them, the quintessence field is one of the most commonly studied candidates. Phantom-like scalar fields, characterized by a negative kinetic term, have also been explored in modeling dark energy Bamba:2012cp ; CALDWELL200223 ; Carroll:2003st ; Copeland:2006wr ; Caldwell:2003vq . Furthermore, several recent works Saha:2024irh ; Debnath:2024urb ; Saha:2024xbg ; Saha:2023zos have investigated models where dynamical dark energy interacts with dark matter, highlighting the growing interest in such interacting scenarios. To exhibit the strength of our proposed solution, we have also worked with dynamical dark energy models and have shown that the equivalence of the coherent oscillation of the ultralight scalar field model and the CDM sector can be maintained properly if we follow the rules of the equivalence as discussed in this paper.
The structure of this paper is as follows. In Section II, we revisit the approach developed by Ratra, outlining the conditions under which a rapidly oscillating scalar field can effectively mimic cold dark matter (CDM) behavior. We also discuss the relevant averaging scheme used in this context. Section III extends this framework to scenarios involving additional cosmic components, allowing for cases where the universe is not purely dust dominated. In Section III.1, we explore how the system evolves when the scalar field is allowed to decay into radiation through a non-minimal coupling. Section III.2 outlines the dependence of the method on initial conditions and provides the initial values and parameter choices used consistently throughout the paper. Section IV presents a detailed numerical analysis based on appropriate initial conditions and the theoretical formulations developed for both minimally and non-minimally coupled cases. Finally, we summarize our results and provide concluding remarks in the last section.
@@ II The traditional model: Only one rapidly oscillating scalar field and CDM-like features
The curvature parameter $k$, taking values 00, $+1$, and $-1$, distinguishes between flat, closed, and open FLRW spacetimes. Here $a(t)$ is the scale factor, which is dimensionless when $k=0$, making the comoving coordinate $r$ carry the dimension of length. For $k=\pm 1$, $a(t)$ has the dimension of length and $r$ becomes dimensionless. In this section, we restrict our analysis to the spatially flat case with $k=0$.
@@
In the above equation, the functions ϕ±subscriptitalic-$\phi_{\pm}$, ϕ˙±subscript˙italic-$\dot{\phi}_{\pm}$, and $H$ vary in the cosmological timescale, whereas the functions $\cos\alpha(t)$, $\sin\alpha(t)$ vary much faster than the aforementioned functions. For this reason, we can assume the terms multiplying $\cos\alpha(t)$ and $\sin\alpha(t)$ in the above equation to be separately equal to zero. This requirement yields:
@@
where the time component of the energy momentum tensor gives, $-T_{0}^{0}=\langle\rho\rangle$ Here, the angular brackets specify the time average over a time period $2\pi/m$. It must be noted that this averaging is over the oscillating time scale of the scalar field, not a complete time average over cosmological time. As a consequence, this time average of various quantities will not eradicate all time dependence of the quantities.

### v2

@@ Abstract
The idea that coherent oscillations of a scalar field, oscillating over a time period that is much shorter than the cosmological timescale, can exhibit cold dark matter (CDM) like behavior was previously established. In our work we first show that this equivalence between the oscillating scalar field model and the CDM sector is exact only in a flat Friedmann-Lemaitre-Robertson-Walker (FLRW) spacetime in the absence of cosmological constant and any other possible matter components in the universe when the mass of the scalar field is very large compared to the Hubble parameter. Then we show how to generalize the equivalence between the coherently oscillating scalar field model and the CDM sector in a spatially curved universe with multiple matter components. Using our general method, we will show how a coherently oscillating scalar field model can represent the CDM sector where it decays into the radiation. Our method is powerful enough to work out the dynamics of gravitational collapse in a closed FLRW spacetime where the coherently oscillating scalar field model represents the CDM sector. We have, for the first time, presented a consistent method which specifies how a coherently oscillating scalar field model, where the scalar field is ultralight, acts like the CDM sector in a multicomponent universe.
@@ I Introduction
[…]
@@
Ratra PhysRevD.44.352 later extended the analysis of Turner and produced a consistent mathematical method to deal with coherently oscillating scalar fields in cosmology. Ratra consistently worked out the background cosmological dynamics of the universe, where the scalar field was the only matter component, and then used a synchronous gauge to calculate cosmological perturbations on the background model. The scalar field dark matter model has been extensively explored in the literature. Several subsequent studies Baer:2014eja ; Hidalgo:2017dfp ; Hotinli:2021vxg ; Garcia:2022vwm ; ChoeJo:2023ffp have explicitly adopted the approach introduced by Turner and Ratra. Hwang HWANG1997241 introduced an alternative approach by employing the uniform-curvature gauge to derive the perturbed equations, demonstrating its suitability for scalar field models. Importantly, although these studies use different gauges for spatial perturbations, the background evolution remains consistent across these approaches. The linear perturbation framework has also been applied to general power-law (scalar field) potentials, with the case $n=2$ treated as a special instance Cembranos:2015oya . This analysis builds upon previous work and demonstrates that the resulting solutions closely resemble the perturbative behavior of dust-like perfect fluids. These methods have been further developed to include second-order perturbations Noh:2013coa and extended to fully nonlinear analysis Noh:2017sdj , reinforcing the result that coherently oscillating scalar fields (or axion-like fields) exhibit cold dark matter-like behavior. Additionally, extensions of the scalar field dark matter framework have considered corrections to the standard behavior McDonald:1993ky ; Cembranos:2018ulm , including higher-order interactions or modifications to the potential. Other studies have incorporated self-interacting terms within the $\phi^{2}$ potential in the context of interacting dark sector models Aboubrahim:2024spa ; vandeBruck:2022xbk , further broadening the scope and applicability of scalar field dark matter scenarios.
@@
Investigations into the dynamical and gravitational instabilities associated with oscillating scalar fields have also played a crucial role in assessing their cosmological viability. A detailed analysis in Ref. Johnson:2008se demonstrated that scalar fields with a negative equation-of-state parameter possess a large-scale dynamical instability for the growth of perturbations. This instability does not occur when the scalar field exhibits a vanishing equation-of-state parameter, making such configurations unsuitable for explaining cosmic acceleration but remaining viable as dark-matter candidates. The growth of cosmological perturbations in scalar field dark matter models has been further examined in works like Alcubierre:2015ipa ; Marsh:2010wq , confirming their relevance for structure formation scenarios. Furthermore, extensive numerical comparisons with the standard cold dark matter paradigm, such as those in Torres:2014bpa , have shown that scalar field and CDM perturbations evolve nearly identically in the linear regime. The primary difference lies in the presence of a characteristic cutoff in the power spectrum of the density perturbation, introduced by the scalar field dynamics. To improve the accuracy of these models, additional effort has yielded the effective fluid approximation for coherently oscillating scalar field dark matter Passaglia:2022bcr , thereby enhancing their consistency with observational data.
From the foundational works on the equivalence of the CDM sector to the scalar field sector PhysRevD.28.1243 ; PhysRevD.44.352 ; HWANG1997241 it can be shown that,
  • the equivalence holds in a spatially flat FLRW spacetime. In the present work, we will show how this restriction of spatial flatness can be addressed. We will generalize the foundational results to a spatially curved FLRW spacetime.
  • In the foundational works, the equivalence only worked when the universe contained a single scalar field. In the present work, we will generalize the equivalence of the CDM sector to a scalar field sector when the universe contains various other matter fields.
  • The foundational works could not address the issue of coupling the scalar field sector to other sectors consistently, as the introduction of other sectors invariably produced a multicomponent universe. In the present work, we will show how one can introduce decaying dark matter in a formally consistent fashion. Turner PhysRevD.28.1243 did work informally with a decaying dark matter model; the present work establishes the formal way such models can be tackled.
  • In the context of interacting models, several studies have adopted the final results from the scalar field dark matter framework developed by Ratra and Turner, applying them directly to systems involving additional components or interactions HWANG20091 ; Mishra:2017ehw ; Park:2012ru ; Cardenas:2020nny ; Brax:2019fzb ; Urena-Lopez:2015gur . In these works, the oscillating scalar field is treated as dark matter without explicitly deriving any conditions under which such generalizations can be done. The authors of Ref. Mishra:2017ehw hinted at the fine-tuning of the initial value of the ultralight scalar field. In the present work, for the first time, we will show that such generalizations require a specific constraint on the initial condition of the system.
  • Our method works for CDM models even in the end phase of radiation domination or in the dark energy-dominated phase. The calculations presented in this paper make the coherently oscillating ultralight scalar field model of CDM very general and applicable to various phases of cosmological dynamics.
@@
The phrase mass of the scalar field frequently appears in the literature, and it is worth clarifying its intended meaning. Typically, it refers to the mass of the quantum of the scalar field—that is, the associated particle. This identification becomes natural and direct when working in natural units, where physical quantities like mass and energy are often interchangeable. In this work, however, we adopt the geometrized unit system in which $8\pi G=c=1$, where $G$ denotes Newton’s gravitational constant and $c$ is the speed of light. Under this convention, both the scalar field mass $m$ and the Hubble parameter $H$ are expressed in units of inverse length, specifically ${\rm cm}^{-1}$. Unlike in natural units, where such quantities would naturally carry dimensions of mass or energy—our chosen system does not assign them such interpretations directly. Furthermore, we do not quantize the scalar field in this work; our analysis is entirely classical. Nevertheless, for ease of communication and consistency with standard terminology in the field, we continue to refer to $m$ as the “mass” of the dark matter particle—while keeping in mind that, in our unit system, it has dimensions of inverse length ${\rm cm}^{-1}$.
[…]
The structure of this paper is as follows. In Section II, we revisit the approach developed by Ratra, outlining the conditions under which a rapidly oscillating scalar field can effectively mimic CDM behavior. We also discuss the relevant averaging scheme used in this context. Section III extends this framework to scenarios involving additional cosmic components, allowing for cases where the universe is not purely dust dominated. In Section III.3, we explore how the system evolves when the scalar field is allowed to decay into radiation. Section III.4 outlines the dependence of the method on initial conditions and provides the initial values and parameter choices used consistently throughout the paper. Section IV presents a detailed numerical analysis based on appropriate initial conditions and the theoretical formulations developed. The next section V gives the results when our method is extended for models involving gravitational collapse. Finally, we summarize our results and provide concluding remarks in the last section.
@@ II The traditional model: Only one rapidly oscillating scalar field and CDM-like features
The curvature parameter $k$, taking values 0, +1+1, and 1-1, distinguishes between flat, closed, and open FLRW spacetimes. Here $a(t)$ is the scale factor, which is dimensionless when $k=0$, making the comoving coordinate $r$ carry the dimension of length. For $k=\pm 1$, $a(t)$ has the dimension of length and $r$ becomes dimensionless. In this section, we restrict our analysis to the spatially flat case with $k=0$.
@@
In the above equation, the functions $\phi_{\pm}$, $\dot{\phi}_{\pm}$, and $H$ vary in the cosmological timescale, whereas the functions $\cos\alpha(t)$, $\sin\alpha(t)$ vary much faster than the aforementioned functions. For this reason, we can assume the terms multiplying $\cos\alpha(t)$ and $\sin\alpha(t)$ in the above equation to be separately equal to zero. This requirement yields:
@@
where the time component of the energy momentum tensor gives, $T_{0}^{0}=\langle\rho\rangle$ . Here, the angular brackets specify the time average over a time period $2\pi/m$. It must be noted that this averaging is over the oscillating time scale of the scalar field, not a complete time average over cosmological time. As a consequence, this time average of various quantities will not eradicate all time dependence of the quantities.
Comparison of Oscillating Scalar Field CDM Models
Model type One scalar field Spatial curvature Multiple components Interacting CDM Special Initial condition
Ratra–Turner type Yes No No No Insensitive
SDB type Yes Yes Yes Yes Sensitive
Table 1: Table showing the comparison of our work with the foundational works on this area. Here “SDB” stands for the last names of authors of the present paper, Saha-Dey-Bhattacharya.
@@
$\langle f(t)\rangle=\frac{m}{2\pi}\int_{0}^{\frac{2\pi}{m}}f(t^{\prime})\,dt^{\prime}\,.$ (15)
Using this averaging scheme we can show that the average energy density of the system is given as:
$\langle\rho\rangle=\frac{1}{2}m^{2}(C_{+}^{2}+C_{-}^{2})\frac{1}{a^{3}}\left(1+\frac{9}{8}\frac{H^{2}}{m^{2}}\right)\,$, (16)
and the average pressure is:
$\langle P\rangle=\frac{9}{16}m^{2}(C_{+}^{2}+C_{-}^{2})\frac{1}{a^{3}}\left(\frac{H^{2}}{m^{2}}\right)\,.$ (17)
The details of the calculation is given in appendix A. The ratio of the pressure and the energy density of the system then yields:
In the limit $m\gg H$, we see that the above ratio tends to zero, specifying an effective equation of state (EoS) of dust. To get the functional form of the scale-factor one has to use the above average values in the first Friedmann equation, Eq. (14). Doing so one obtains:
$a(t)=a_{i}\left[1+A(t-t_{i})\right]^{2/3}\,$, (19)
where $a_{i}$ is the scale-factor at $t=t_{i}$ and
$A^{2}\equiv\frac{3m^{2}}{8a_{i}^{3}}(C_{+}^{2}+C_{-}^{2})\,.$ (20)
This is the matter-dominated universe solution. It is shown in appendix A that when $H^{2}\gg m^{2}$ and $\langle P\rangle\ll\langle\rho\rangle$, the other Friedmann equation (the acceleration equation) is also satisfied. The crucial point about this solution is that only the above matter-dominated solution of the scale-factor satisfies Eq. (12). It is seen from the above calculations when there is only one coherently oscillating scalar field in the universe, oscillating in a time scale which is much shorter than the cosmological time scale, the cosmological system behaves like a dark matter dominated system and only in these case both Eq. (11) and Eq. (12) are simultaneously satisfied. If the scale-factor is different from the form of the scale-factor in pure matter domination, then Eq. (12) will never be satisfied and the whole calculational scheme will become inconsistent.
In the next section we will show how this model can be generalized for a multicomponent universe which can be even spatially curved. Our generalized model for oscillating ultralight scalar fields can even accommodate interaction with other components of the universe. The Ratra-Turner model is insensitive to initial condition as the equivalence holds only when the scalar field solution is given by the expression in Eq. (13) and we do not have to specify some specific initial condition for the scalar field. We do not obtain Eq. (13) for some particular initial condition. The generalization of the Ratra-Turner like model requires a particular kind of initial condition of the problem. This topic will be discussed in the next section. A brief comparison of the various oscillating ultralight scalar field models, producing CDM like features, is presented in Table 1.
  1. 1.
    The separation of mass or time scales, given by $m\gg H$.
  2. 2.
    The energy density $\rho$ is solely coming from a dust solution.
$=0\,$,
$[\ddot{\phi}_{+}-2m\dot{\phi}_{-}+3H(\dot{\phi}_{+}-m\phi_{-})]\sin\alpha$
$\ddot{\phi}_{+}-2m\dot{\phi}_{-}+3H(\dot{\phi}_{+}-m\phi_{-})$ $=$ $0\,$, (22)
$\ddot{\phi}_{-}+2m\dot{\phi}_{+}+3H(\dot{\phi}_{-}+m\phi_{+})$ $=$ $0\,.$ (23)
If we again segregate the terms in the above equation depending upon the powers of $m$ then we will get back Eq. (11) and Eq. (12), which will only produce a purely dust dominated universe. This point was explained in the first paragraph of this section. Consequently, in the presence of other matter sources, we do not simplify the above solutions and work with Eq. (22) and Eq. (23) as a set of coupled differential equations in $\phi_{\pm}$. Here we assume that the universe is not purely dust dominated, although $m\gg H$. In the presence of another perfect fluid component and in the presence of spatial curvature, the Friedmann equations are: